### v3

@@ A.3 Distributional Alignment
Define $\mathcal{L}(\boldsymbol{w})\leftarrow\text{WassersteinDistance}(\text{EstimateDistribution}(\mathcal{X},\boldsymbol{w}),q)$.
while a valid weight vector $\boldsymbol{w}$ is not found do
[…]
@@
As illustrated in Figure 9(a), an observed behavior distribution can be recreated from a mixture of behavioral codes. Algorithm 2 outlines the procedure for determining the optimal mixture of codes to achieve distributional alignment with an observed distribution.
After optimizing the weights ($\boldsymbol{w}$) over behavioral codes ($\mathcal{X}$) as described in Algorithm 2, the mixture of codes is evaluated by comparing the LLM-elicited behavior distribution with the observed human behavior distribution, as shown in Figure 9(b). For this evaluation, the mixture generates 1000 samples by sampling a behavioral code according to the weights and eliciting a behavior.
@@
Table 3: Distributional alignment performance is evaluated using the Wasserstein metric, which measures the distance between the LLM-elicited behavior distributions (with or without the behavior code mixture) or a subsample human distribution and the observed human distribution for each game. A lower Wasserstein distance indicates a closer alignment. “” denotes distributions that are statistically indistinguishable from the observed human distribution under the Kolmogorov–Smirnov (KS) test ($p>0.05$), while “” indicates indistinguishability under a relaxed KS test, where behavior choices are rounded to multiples of 5 (multiples of 2 for Public Goods) before statistical testing. The results suggest that using a mixture of system prompts improves the alignment of LLM behavior distributions with human behavior distributions compared to the default system prompt, highlighting AI’s ability to decipher behavior distributions.
The distributional alignment successfully fits the behavioral codes to a distribution of human play. Table 3 shows that the LLM-elicited distribution given the deciphered mixture of behavioral codes is indistinguishable from the human distribution in 5 out of 7 games, under a relaxed Kolmogorov-Smirnov (KS) test [^2] In the relaxed Kolmogorov-Smirnov (KS) test, samples are binned with a width of 10 (the bin width for the Public Goods game is 2)..
Table 3 presents the quantitative evaluation results. To provide context, two baselines are included: (1) An upper bound: 1000 random samples are drawn from the human-play records to compare against the full human behavior distribution. (2) A lower bound: 1000 behavior choices are generated using the default system prompt without leveraging the behavioral codes. The evaluation results demonstrate that using a mixture of system prompts improves the alignment of LLM behavior distributions with human behavior distributions compared to the default system prompt, highlighting AI’s ability to decipher and then elicit behavior distributions.
Appendix B Interpretability of Behavioral Codes
To verify the interpretability of the learned behavioral codes, we perform a textual analysis. Specifically, each behavioral code is represented as a bag of words, and the TF-IDF importance score of each word within the code is calculated. Table 4 presents the top 50 keywords from the behavioral codes for each game scenario.
@@ Appendix B Interpretability of Behavioral Codes
decision proposal decision decision investor group risk
fairness decision fairness risk profit contribution decision
strategic player proposal investment decision benefit reward
maker strategic ensure potential ensure decision potential
aim ensure outcome return balance personal high
balance party benefit aim term collective maximize
outcome aim reflect balance strategic aim balance
benefit outcome strategic strategic trust maximize gain
ensure offer maker reflect maximize balance aim
choice fairness offer investor aim contribute outcome
reflect maximize aim maximize future individual choice
generosity acceptance fair approach benefit strategic probability
party benefit value choice gain outcome maker
maximize accept gain growth long payoff point
fair balance maximize make relationship ensure strategic
value fair balance ensure maintain gain approach
make likely accept high return overall minimize
consider consider prioritize outcome foster make make
maintain propose consider gain fairness return carefully
prioritize maker high conservative fair maker achieve
self mutual secure maker banker high optimal
resource negotiator evaluate cautious investment resource strategist
advantage make rational trust encourage consider loss
thoughtful goal reasonable maintain party success calculate
slightly generosity share moderate reflect prioritize strategy
generous likelihood equitable reward cooperation reflect optimize
create focus achieve carefully player optimize safety
gain understand choice resource maker participant scenario
positive generous substantial achieve value cooperative avoid
sense reflect focus seek mutual approach ensure
goodwill gain receive balanced immediate term prioritize
strive cooperation strong strategy outcome focus excel
favor create scenario prioritize consider carefully reflect
approach value carefully optimize ongoing game possible
focus prioritize self consider focus long cautious
reasonable foster possible prudent interaction potential conservative
considerate positive standard significant continue strategy significant
scenario high sense invest collaboration optimal consider
equitable highly pragmatic minimize prioritize goal involve
optimize beneficial negotiation calculate optimize scenario focus
foster understanding acceptable prefer importance possible evaluate
involve reasonable negotiator opportunity positive effort favor
seek importance personal thoughtful goal enhance goal
share empathetic goal careful understand cooperation level
balanced optimize significant level self collaborative choose
personal involve threshold substantial secure achieve option
retain achieve make calculated reward balanced success
achieve self meet player strive foster calculated
understand agreement maintain caution feel slightly thinker
offer perceive advantageous small win cautious bold
@@
Table 4: Top 50 keywords in behavioral codes for each game scenario. Keywords are sorted by TF-IDF word importance.
@@
To understand which keywords influence the behaviors, as deciphered by LLMs, we perform a linear regression analysis. Treating each behavioral code as an observation, we generate 10 behavior choices with this code and predict the mean behavior based on the occurrence of keywords. An ordinary least squares (OLS) linear regression model is applied using one-hot keyword occurrence feature vectors with 50 dimensions. Table 5 presents the complete linear regression results, showing the impact of each keyword in the game scenarios. Table 1 highlights the top 10 keywords in the linear regression analysis.
@@
Table 5: Linear regression analysis of behaviors based on keywords in the behavioral codes. Specifically, each behavioral code is transformed into a one-hot keyword-occurrence feature vector of 50 dimensions to predict the mean behavior that can be elicited by that code. This prediction is performed using an (OLS) linear regression model. The results demonstrate that keyword occurrences in behavioral codes effectively predict the behaviors of LLMs, offering interpretability of the behavioral codes.
@@
We further explore the deciphered information contained in the obtained behavioral codes by analyzing the correlations between the elicited behaviors and the principal components (PCs) of the behavioral codes. The PCs are derived from the 50-dimensional one-hot keyword vectors representing the codes for each game. Each code is used to elicit a behavior 10 times, and the Spearman’s correlation between the mean elicited behavior and the PC value is computed. Figure 10 shows the first five PCs, where moderate to strong correlations are observed, demonstrating the effectiveness of behavioral codes in deciphering behaviors.
Figure 10: Correlations between the elicited behaviors and behavioral codes’ principal components (PCs). Principal components are obtained from the 50-dimensional, one-hot keyword vectors of the codes of each game. Spearman’s correlation is calculated between the PC score of each behavioral code and the mean of elicited behaviors. The figure displays the first five PCs.
@@
The interpretability of behavioral codes is also consistent and generalizable. As illustrated in Fig 11, the smoothness of behaviors regarding the behavioral code semantics can be observed: Behavioral codes that are close in the semantic space tend to elicit similar behaviors, highlighting the structured and predictable nature of their influence on behavior generation.
@@
Figure 11: The 2D projection of the behavioral codes across all games. For each behavioral code, the mean behavior that is elicited is calculated (and normalized to the range $[0,1]$). Corresponding colors are applied to the codes based on these values. This visualization highlights the continuity of the behavioral codes with respect to their elicited behaviors.
Appendix C Game Characteristics and Correlations
@@
Behavioral codes offer new insights into games. From the keywords used in game scenarios (Table 4), we observe notable inter-game relationships. To further dissect these correlations, we analyze behavioral codes in a semantic space. Specifically, each behavioral code is embedded into a 1536-dimensional space using the embedding model OpenAI Ada 3 (text-embedding-3-large). These high-dimensional vectors are projected onto a 2D low-dimensional map using UMAP. The resulting projection map is displayed in Figure 4, with a heatmap highlighting semantic similarities between games. The similarities are computed as the aggregated cosine similarity of the high-dimensional embeddings.
Appendix D Population Signatures
Figure 12: Behavior distribution of different populations under the Dictator game.
@@ Appendix D Population Signatures
By aligning the behavior distribution with a target population, as detailed in Sec. A.3, LLMs effectively decipher the signatures of this population into a mixture of behavioral codes.
@@
Behavioral codes can also serve as a valuable tool for identifying distinct decision-making patterns across different populations. In a meta-study of the Dictator Game [9] [^3] Data from Engel [9]: https://osf.io/xc73h/, retrieved on March 14, 2025., various populations were analyzed, revealing distinct behavioral distributions. Figure 12 displays the behavior distributions of various populations in the Dictator game, including comparisons between Students and Non-students populations, as well as populations from diverse societal backgrounds. Figure 6 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies.

### v5

@@ 4.3 Distributional Alignment
Define $\mathcal{L}(\boldsymbol{w})\leftarrow\text{WassersteinDistance}(\text{EstimateDistribution}(\mathcal{X},\boldsymbol{w}),q)+\alpha\cdot\|\boldsymbol{w}\|$
while a valid weight vector $\boldsymbol{w}$ (i.e., $\sum w_{i}=1$ and $w_{i}\in[0,1]$) is not found do
     $\boldsymbol{w}\leftarrow$ a random vector in which each element is uniformly sampled from [0,1) and then normalized to sum up to 1
@@
As illustrated in Figure 3(a), an observed behavior distribution can be recreated from a mixture of behavioral codes. Algorithm S2 outlines the procedure for determining the optimal mixture of codes to achieve distributional alignment with an observed distribution. For the OptimizeVector function in our experiments, we used the Sequential Least Squares Programming (SLSQP) algorithm from the scikit-learn package. The optimization was performed with the linear constraint $\sum w_{i}=1$ and bounds $w_{i}\in[0,1]$. We set the termination tolerance to $10^{-6}$ and the regularization parameter $\alpha$ to 3.
Figure S3: Illustration of how LLM can decipher and elicit behavior distributions.
After optimizing the weights ($\boldsymbol{w}$) over behavioral codes ($\mathcal{X}$) as described in Algorithm S2, the mixture of codes is evaluated by comparing the LLM-elicited behavior distribution with the observed human behavior distribution, as shown in Figure 3(b). For this evaluation, the mixture generates 1000 samples by sampling a behavioral code according to the weights and eliciting a behavior.
The distributional alignment successfully fits the behavioral codes to a distribution of human play. Table S2 shows that the LLM-elicited distribution given the deciphered mixture of behavioral codes is indistinguishable from the human distribution in 5 out of 7 games, under a relaxed Kolmogorov-Smirnov (KS) test [^2] In the relaxed Kolmogorov-Smirnov (KS) test, samples are binned with a width of 10 (the bin width for the Public Goods game is 2)..
Table S2 also presents the distribution similarity. To provide context, two baselines are included:
  1. (1)
    An upper bound: 1000 random samples are drawn from the human-play records to compare against the full human behavior distribution.
  2. (2)
    A lower bound: 1000 behavior choices are generated using the default system prompt (“You are a helpful assistant.”) without leveraging the behavioral codes.
The evaluation results demonstrate that using a mixture of system prompts improves the alignment of LLM behavior distributions with human behavior distributions compared to the default system prompt, highlighting AI’s ability to decipher and then elicit behavior distributions.
Table S2: Distributional alignment performance. Columns represent game scenarios, and rows indicate different behavior generation/sampling methods. The elicited/sampled behaviors are then compared with the ground-truth human distributions. The performance is evaluated using the Wasserstein metric, which measures the distance between the LLM-elicited behavior distributions (with or without the behavior code mixture) or a subsample human distribution and the observed human distribution for each game. A lower Wasserstein distance indicates a closer alignment. “” denotes distributions that are statistically indistinguishable from the observed human distribution under the Kolmogorov–Smirnov (KS) test ($p>0.05$), while “” indicates indistinguishability under a relaxed KS test, where behavior choices are rounded to multiples of 5 (multiples of 2 for Public Goods) before statistical testing. The results suggest that using a mixture of behavioral codes improves the alignment of LLM behavior distributions with human behavior distributions compared to the default system prompt, highlighting AI’s ability to decipher behavior distributions.
@@
We also investigate how learned behavioral code mixtures can generalize across games. We elicit behaviors from one (target) game using behavioral codes from another (source) game, and compared outcomes with human-play records in the target game. The Wasserstein distances of the elicited behavior distribution and the human distribution are shown in Table S3. In particular, we see that the behavioral codes obtained from one game can elicit behavior distributions that are closer to the human distribution than the default system prompt (no behavioral code) in most game settings, where the most aligned behaviors are still elicited from codes derived from the same game.
Table S3: The performance of using the behavioral code mixture learned from one game (source) to another game (target). Columns represent target game scenarios, and rows indicate the source games where the behavioral code mixtures are learned from. The “Default System Prompt“ stands for using the default system prompt without leveraging the behavioral codes. The elicited behaviors are then compared with the ground-truth human distributions. The performance is evaluated using the Wasserstein metric, which measures the distance between the LLM-elicited behavior distributions (with or without the behavior code mixture) or a subsample human distribution and the observed human distribution for each game. A lower Wasserstein distance indicates a closer alignment.
@@
Default System Prompt 25.58 11.95 32.53 42.78 20.35 24.00 26.99
Dictator 0.74 17.94 12.77 8.49 4.05 20.83 9.49
Proposer 22.41 0.92 9.37 15.29 19.86 8.11 10.10
Responder 11.89 10.63 1.11 25.11 14.73 13.78 11.33
Investor 17.99 9.38 14.86 2.30 25.09 8.53 22.41
Banker 5.67 17.57 18.73 13.37 1.52 24.88 12.19
Public Goods 15.16 7.28 15.04 9.60 20.70 2.27 13.35
Bomb 16.53 8.89 5.63 19.25 20.61 13.68 4.78
5 Supplementary Text
5.1 Interpretability of Behavioral Codes
To verify the interpretability of the learned behavioral codes, we perform a textual analysis. Specifically, we first preprocess the behavioral codes using the spaCy library with the en_core_web_sm model. This preprocessing involved tokenization, lemmatization, and the removal of stop words (using .is_stop) and non-alphabetic tokens (using .is_alpha). Each preprocessed behavioral code is then represented as a bag of keywords, and the TF-IDF importance score of each keyword within the code is calculated. Table S4 presents the top 50 keywords (those with highest aggregated TF-IDF scores) from the behavioral codes for each game scenario.
@@ 5.1 Interpretability of Behavioral Codes
Table S4: Top 50 keywords in behavioral codes for each game scenario. Each column represent a game scenario, and keywords are list in rows, sorted by TF-IDF word importance.
Dictator Proposer Responder Investor Banker Public Goods Bomb
1 decision proposal decision decision investor group risk
2 fairness decision fairness risk profit contribution decision
3 strategic player proposal investment decision benefit reward
4 maker strategic ensure potential ensure decision potential
5 aim ensure outcome return balance personal high
6 balance party benefit aim term collective maximize
7 outcome aim reflect balance strategic aim balance
8 benefit outcome strategic strategic trust maximize gain
9 ensure offer maker reflect maximize balance aim
10 choice fairness offer investor aim contribute outcome
11 reflect maximize aim maximize future individual choice
12 generosity acceptance fair approach benefit strategic probability
13 party benefit value choice gain outcome maker
14 maximize accept gain growth long payoff point
15 fair balance maximize make relationship ensure strategic
16 value fair balance ensure maintain gain approach
17 make likely accept high return overall minimize
18 consider consider prioritize outcome foster make make
19 maintain propose consider gain fairness return carefully
20 prioritize maker high conservative fair maker achieve
21 self mutual secure maker banker high optimal
22 resource negotiator evaluate cautious investment resource strategist
23 advantage make rational trust encourage consider loss
24 thoughtful goal reasonable maintain party success calculate
25 slightly generosity share moderate reflect prioritize strategy
26 generous likelihood equitable reward cooperation reflect optimize
27 create focus achieve carefully player optimize safety
28 gain understand choice resource maker participant scenario
29 positive generous substantial achieve value cooperative avoid
30 sense reflect focus seek mutual approach ensure
31 goodwill gain receive balanced immediate term prioritize
32 strive cooperation strong strategy outcome focus excel
33 favor create scenario prioritize consider carefully reflect
34 approach value carefully optimize ongoing game possible
35 focus prioritize self consider focus long cautious
36 reasonable foster possible prudent interaction potential conservative
37 considerate positive standard significant continue strategy significant
38 scenario high sense invest collaboration optimal consider
39 equitable highly pragmatic minimize prioritize goal involve
40 optimize beneficial negotiation calculate optimize scenario focus
41 foster understanding acceptable prefer importance possible evaluate
42 involve reasonable negotiator opportunity positive effort favor
43 seek importance personal thoughtful goal enhance goal
44 share empathetic goal careful understand cooperation level
45 balanced optimize significant level self collaborative choose
46 personal involve threshold substantial secure achieve option
47 retain achieve make calculated reward balanced success
48 achieve self meet player strive foster calculated
49 understand agreement maintain caution feel slightly thinker
50 offer perceive advantageous small win cautious bold
@@
To understand which keywords influence the behaviors, as deciphered by LLMs, we perform a linear regression analysis. Treating each behavioral code as an observation, we generate 10 behavior choices with this code and predict the mean behavior based on the occurrence of keywords. An ordinary least squares (OLS) linear regression model is applied using binary keyword occurrence feature vectors with 50 dimensions. Table S5 presents the complete linear regression results, showing the impact of each keyword in the game scenarios. Figure 3 highlights the top 10 keywords in the linear regression analysis. To alleviate the influence of multicollinearity, we also report the LASSO regression results in Table S6. A comparison of the OLS and LASSO results reveals strong alignment among the top keywords (those with high absolute LR coefficients), suggesting the robustness of the observed correlations.
@@
Table S5: Linear regression analysis of behaviors based on keywords in the behavioral codes. Specifically, each behavioral code is transformed into a binary keyword-occurrence feature vector of 50 dimensions to predict the mean behavior that can be elicited by that code. This prediction is performed using an ordinary least squares (OLS) linear regression model. The results demonstrate that keyword occurrences in behavioral codes effectively predict the behaviors of LLMs, offering interpretability of the behavioral codes.
Table S6: Linear regression analysis of behaviors based on keywords in the behavioral codes. Specifically, each behavioral code is transformed into a binary keyword-occurrence feature vector of 50 dimensions to predict the mean behavior that can be elicited by that code. This prediction is performed using a LASSO linear regression model with the regularization coefficient ($\alpha=0.3$). The results demonstrate that keyword occurrences in behavioral codes effectively predict the behaviors of LLMs, offering interpretability of the behavioral codes.
@@
We further explore the deciphered information contained in the obtained behavioral codes by analyzing the correlations between the elicited behaviors and the principal components (PCs) of the behavioral codes. The PCs are derived from the 50-dimensional binary keyword vectors representing the codes for each game. Each code is used to elicit a behavior 10 times, and the Spearman’s correlation between the mean elicited behavior and the PC value is computed. Figures S4-S10 show the first five PCs and their loading vectors, where moderate to strong correlations are observed, demonstrating the effectiveness of behavioral codes in deciphering behaviors. Figure S11 show the decay of the explained variance ratios for the PCs of keywords in each game. The ratios decay smoothly, aligning with the common trend observed when applying SVD to sparse document-keyword vectors  (?).
Figure S4: Correlations between the elicited behaviors and behavioral codes’ PCs (left) and the loading vectors of PCs (right) for the Dictator game.
Figure S5: Correlations between the elicited behaviors and behavioral codes’ PCs (left) and the loading vectors of PCs (right) for the Proposer game.
Figure S6: Correlations between the elicited behaviors and behavioral codes’ PCs (left) and the loading vectors of PCs (right) for the Responder game.
Figure S7: Correlations between the elicited behaviors and behavioral codes’ PCs (left) and the loading vectors of PCs (right) for the Investor game.
Figure S8: Correlations between the elicited behaviors and behavioral codes’ PCs (left) and the loading vectors of PCs (right) for the Banker game.
Figure S9: Correlations between the elicited behaviors and behavioral codes’ PCs (left) and the loading vectors of PCs (right) for the Public Goods game.
Figure S10: Correlations between the elicited behaviors and behavioral codes’ PCs (left) and the loading vectors of PCs (right) for the Bomb game.
Figure S11: Explained variance ratio for each principle component (PC). The x-axis indicates the 50 PCs in order, and the y-axis is the explained variance ratio.
@@
The interpretability of behavioral codes is also consistent and generalizable. As illustrated in Fig S12, the smoothness of behaviors regarding the behavioral code semantics can be observed: Behavioral codes that are close in the semantic space tend to elicit similar behaviors, highlighting the structured and predictable nature of their influence on behavior generation.
@@
Figure S12: The 2D projection of the behavioral codes across all games. For each behavioral code, the mean behavior that is elicited is calculated (and normalized to the range $[0,1]$). Corresponding colors are applied to the codes based on these values. This visualization highlights the continuity of the behavioral codes with respect to their elicited behaviors.
5.2 Game Characteristics and Correlations
@@
Behavioral codes offer new insights into games. From the keywords used in game scenarios (Table S4), we observe notable inter-game relationships. To further dissect these correlations, we analyze behavioral codes in a semantic space. Specifically, each behavioral code is embedded into a 1536-dimensional space using the embedding model OpenAI Ada 3 (text-embedding-3-large). These high-dimensional vectors are projected onto a 2D low-dimensional map using UMAP. The resulting projection map is displayed in Figure 5, with a heatmap highlighting semantic similarities between games. The similarities are computed as the aggregated cosine similarity of the high-dimensional embeddings.
5.3 Behavioral Signatures for Populations
@@
By aligning the behavior distribution with a target population, as detailed in Sec. 4.3, LLMs effectively decipher the signatures of this population into a mixture of behavioral codes.
@@
Figure S13 presents an alternative annotation for the layout, derived through topic modeling. We apply the LDA topic modeling algorithm to the behavioral codes, with each code represented as a 50-dimensional binary keyword vector. A high alignment is observed between the semantic clusters in Figure 6 and the topic modeling results, yielding a Normalized Mutual Information (NMI) score of 0.594 between the clusters and the dominant topics of the codes.
Our topic modeling approach employed the Latent Dirichlet Allocation (LDA) class within the scikit-learn package. We set the number of topics (n_components) to 20 and used a random state seed of 42 for reproducibility. Prior to modeling, behavioral codes were preprocessed using the spaCy library (specifically, the en_core_web_sm model). This preprocessing involved tokenization, lemmatization, and the removal of stop words (using .is_stop) and non-alphabetic tokens (using .is_alpha). Subsequently, these processed codes were vectorized using scikit-learn’s CountVectorizer, configured to remove ’english’ stop words and limit features to a maximum of 1000.
Figure S13: Topic modeling results of the behavioral codes shown in the 2D projection map. We apply the LDA topic modeling algorithm to the behavioral codes, each represented as a 50-dimensional binary keyword vector. Behavioral codes are colored by their dominant topic. Topics are summarized by their most correlated keywords.
Behavioral codes can also serve as a valuable tool for identifying distinct decision-making patterns across different populations. In a meta-study of the Dictator Game  (?) [^3] Data from  (?): https://osf.io/xc73h/, retrieved on March 14, 2025., various populations were analyzed, revealing distinct behavioral distributions. Figure S14 displays the behavior distributions of various populations in the Dictator game, including comparisons between Students and Non-students populations, as well as populations from diverse societal backgrounds. Figure 7 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies.
Figure S14: Behavior distribution of different populations under the Dictator game.